Derby Company’s budgeted sales and direct materials purchases are as follows.

Budgeted Sales

Budgeted Purchases

January

$190,000

$30,000

February

210,000

35,000

March

300,000

45,000

Derby’s sales are collected 10% in the month of sale, 50% in the month following sale, and 36% in the second month following sale; 4% are uncollectible.

Derby’s purchases are paid 60% in the month of purchase, and 40% in the month following purchase.

A. Prepare a schedule of expected collections from customers for the first quarter.

January February March Quarter
January
February
March

Total payments

B. Prepare a schedule of expected payments for direct materials for the first quarter.

January February March Quarter
January
February
March

Total payments
Answer #1

a) Schedule of expected cash collection

January February march Quarter
January 190000*10% = 19000 190000*50% = 95000 190000*36% = 68400 182400
February 210000*10% = 21000 210000*50% = 105000 126000
march 300000*10% = 30000 30000
Total collection 19000 116000 203400 338400

b) Prepare expected payment for direct material

January February March Quater
January 30000*60% = 18000 30000*40% = 12000 30000
February 35000*60% = 21000 35000*40% = 14000 35000
March 45000*60% = 27000 27000
Total payment 18000 33000 41000 92000